• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EDT 05:54
CEST 11:54
KST 18:54
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Finals Preview: Sweet Victory!9[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5
Community News
[$1250] International ShieldBattery League14Weekly Cups (October 5-11): Farewell to 5.0.162Rollback Netcode in Brood War1185.0.17 Patch Notes - October 7, 202624Weekly Cups Results (Sep 28-Oct 4)0
StarCraft 2
General
Server Blocker Terran AddOns placement SCII needs to put more weight on strategy and tact How do you feel about the mass reverts in the 5.0.17 PTR? Team Liquid Map Contest #22: Results and Winners
Tourneys
Sparkling Tuna Cup - Weekly Open Tournament Master Swan Open (Global Bronze-Master 2) $3,500 WardiTV Not-Summer Championship Stellar Fest TWO the Moon (Dec 16-20) ISSL (IntoTheTV x SOOP SC2 League): Premier
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
[G] How to playtest arcade maps [A] Nemrods 1/4 players [M] (2) Sweltering Sands [M] (2) Frigid Storage
External Content
The PondCast: SC2 News & Results Mutation # 547 Temple Of Terror Mutation # 546 Catch the Train Mutation # 545 And Drops and Rifts
Brood War
General
Rollback Netcode in Brood War [Megathread] Daily Funny Videos [Recap] Hive ZvZ in MonstarZ vs Dreams [ASL22] Finals Preview: Sweet Victory! Bot on ladder
Tourneys
[ASL22] Grand Finals [$1250] International ShieldBattery League 2v2v2v2 Tournament Small VOD Thread 2.0
Strategy
Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers Replay Review Process - What do you do?
Other Games
General Games
Stormgate/Frost Giant Megathread Nintendo Switch Thread Dawn of War IV General RTS Discussion Thread Warcraft III: The Frozen Throne
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
US Politics Mega-thread Russo-Ukrainian War Thread Artificial Intelligence Thread The Chess Thread Things Aren’t Peaceful in Palestine
Fan Clubs
Serral Fan Club
Media & Entertainment
[Manga] One Piece Movie Discussion! Diablo Animated Series on Netflix
Sports
MLB/Baseball 2023 Football (Soccer) Thread TeamLiquid Health and Fitness Initiative For 2023
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
The Link Between Gaming and …
TrAiDoS
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
38 yo Retired SWE loo…
PurE)Rabbit-SF
Can Bots Beat Pros?? Starcr…
namkraft
[meme] I finally understa…
LUCKY_NOOB
Customize Sidebar...

Website Feedback

Closed Threads



Active: 9565 users

Mathematics (again) - Page 2

Blogs > achristes
Post a Reply
Prev 1 2 3 Next All
achristes
Profile Blog Joined March 2011
Norway653 Posts
February 01 2012 15:25 GMT
#21
On February 02 2012 00:20 fabiano wrote:
I found:

1 - 0,195 = 0.805 = 80,5%

2007 -> 225.400 - (225.400 x 0,195) = 181.447 NOK
2006 -> 280.000 - (280.000 x 0,195) = 225.400 NOK
2005 -> 280.000 NOK
2004 -> (280.000 x 0,195) + 280.000 = 334.600 NOK
2003 -> (334.600 x 0,195) + 334.600 = 399.847 NOK

1.a) 80,5%
1.b) 181.477 NOK
1.c) 399.847 NOK

What did I do wrong? oO

Switch the bolded ones with 0.805 and try again.

Also, 1a = 0.805, yes you could write it like you did, but it's not answering the question.
youtube.com/spooderm4n | twitch.tv/spooderm4n | Random videos and games I feel like uploading
achristes
Profile Blog Joined March 2011
Norway653 Posts
February 01 2012 15:31 GMT
#22
On February 02 2012 00:24 DarkPlasmaBall wrote:
OP, your 2a is wrong as I explained before. Conditional probability doesn't work in that way because the two events are independent. The second event isn't conditioned on the first.

It's like saying, "I have a six-sided die and a coin. If I roll a 4 first, what's the chance of me flipping heads?" It's still 1/2. It doesn't change just because you roll the die in a certain way.

You didn't ask what the chance of Tor not going to a meeting *and* Nadia going to meeting is simultaneously (in that case, you would carry out the calculation the same way you did). The wording (particularly the "If") makes it a different question.

Nadia's probability of going to a meeting is independent of Tor, so whether or not Tor goes is irrelevant when deciding if Nadia goes. Therefore, Nadia's chances of going is still the established 70%. You don't need to include Tor's chance of not going, because- as was explicitly written in the instructions- the two events (Tor going and Nadia going) are independent events.

If it was slightly reworded as "What's the chance of Tor not going *and* Nadia going to the same meeting", then it would be .4 * .3 = .12.

Hope that helps

Oops, I'm so bad at writing these kinds of questions in english, thanks for telling me

But the bolded part should be .7 because it's the probability of Nadia going is 70%=0.7, or am I being stupid now?
youtube.com/spooderm4n | twitch.tv/spooderm4n | Random videos and games I feel like uploading
Treehead
Profile Blog Joined November 2010
999 Posts
February 01 2012 15:31 GMT
#23
On February 02 2012 00:03 achristes wrote:
Show nested quote +
On February 01 2012 23:39 Caller wrote:
This isn't math this is arithmetic

That statement makes no sense to me, because if you use numbers to find other number I call it math.


Mathematicians are constantly delving into supremely abstract realms to try to discover new theorems, corrollaries, etc. (mathematician's statements of truth) about extremely complicated situations. Some theorists work in areas that realistically may only be useful in practical ways to certain branches of physics. Mathematicians are creative thinkers, critical thinkers and problem solvers.

The OP is primarily concerned with questions of interpretation and execution of equations. Don't see the difference?

A problem for a mathematician goes something more like this:

The 'flea and comb space' is a topological space defined by a subspace of the two-dimensional coordinate plane which contains the point (0,1), all points (x,0), and all points (1/n,y), where n is any positive integer, and x and y are any real numbers between 0 and 1. Additionally, say that a point is "dense" (calling it "dense" because I actually don't remember the real term off hand) in the flea and comb set if it has points which are in the flea and comb set which are infinitely close to it, i.e. for any given arbitrarily small number, there is a point within that distance whcih is in the set. Prove that every point in this space is "dense" in the space, but that it is not the case that every point which is "dense" in the set is necessarily in the set.

The above is a relatively simple and straightforward math problem. More complicated problems may require a day or more of thought and reflection, experimentation and failure to resolve. This is the reason mathematicians bristle at the idea that saying 70% of $50 is $35 is "math" in the same sense. Math requires much more thought and effort.
DarkPlasmaBall
Profile Blog Joined March 2010
United States46537 Posts
Last Edited: 2012-02-01 15:36:49
February 01 2012 15:33 GMT
#24
On February 02 2012 00:31 achristes wrote:
Show nested quote +
On February 02 2012 00:24 DarkPlasmaBall wrote:
OP, your 2a is wrong as I explained before. Conditional probability doesn't work in that way because the two events are independent. The second event isn't conditioned on the first.

It's like saying, "I have a six-sided die and a coin. If I roll a 4 first, what's the chance of me flipping heads?" It's still 1/2. It doesn't change just because you roll the die in a certain way.

You didn't ask what the chance of Tor not going to a meeting *and* Nadia going to meeting is simultaneously (in that case, you would carry out the calculation the same way you did). The wording (particularly the "If") makes it a different question.

Nadia's probability of going to a meeting is independent of Tor, so whether or not Tor goes is irrelevant when deciding if Nadia goes. Therefore, Nadia's chances of going is still the established 70%. You don't need to include Tor's chance of not going, because- as was explicitly written in the instructions- the two events (Tor going and Nadia going) are independent events.

If it was slightly reworded as "What's the chance of Tor not going *and* Nadia going to the same meeting", then it would be .4 * .3 = .12.

Hope that helps

Oops, I'm so bad at writing these kinds of questions in english, thanks for telling me

But the bolded part should be .7 because it's the probability of Nadia going is 70%=0.7, or am I being stupid now?


I messed up the ending part to my previous post (should be ".4 * .7 = .28.") Edited it ^^

.4 = 1-.6 is P(Tor not going)
.7 is P(Nadia going)
If it were worded differently.
"There is nothing more satisfying than looking at a crowd of people and helping them get what I love." ~Day[9] Daily #100
achristes
Profile Blog Joined March 2011
Norway653 Posts
February 01 2012 15:36 GMT
#25
On February 02 2012 00:31 Treehead wrote:
Show nested quote +
On February 02 2012 00:03 achristes wrote:
On February 01 2012 23:39 Caller wrote:
This isn't math this is arithmetic

That statement makes no sense to me, because if you use numbers to find other number I call it math.


Mathematicians are constantly delving into supremely abstract realms to try to discover new theorems, corrollaries, etc. (mathematician's statements of truth) about extremely complicated situations. Some theorists work in areas that realistically may only be useful in practical ways to certain branches of physics. Mathematicians are creative thinkers, critical thinkers and problem solvers.

The OP is primarily concerned with questions of interpretation and execution of equations. Don't see the difference?

A problem for a mathematician goes something more like this:

The 'flea and comb space' is a topological space defined by a subspace of the two-dimensional coordinate plane which contains the point (0,1), all points (x,0), and all points (1/n,y), where n is any positive integer, and x and y are any real numbers between 0 and 1. Additionally, say that a point is "dense" (calling it "dense" because I actually don't remember the real term off hand) in the flea and comb set if it has points which are in the flea and comb set which are infinitely close to it, i.e. for any given arbitrarily small number, there is a point within that distance whcih is in the set. Prove that every point in this space is "dense" in the space, but that it is not the case that every point which is "dense" in the set is necessarily in the set.

The above is a relatively simple and straightforward math problem. More complicated problems may require a day or more of thought and reflection, experimentation and failure to resolve. This is the reason mathematicians bristle at the idea that saying 70% of $50 is $35 is "math" in the same sense. Math requires much more thought and effort.

So what you are saying is that mathematicians don't like that "normal" people say that the simpler parts of math is math?
I can understand it if that is the case though, but what am I supposed to call it then? ^^
youtube.com/spooderm4n | twitch.tv/spooderm4n | Random videos and games I feel like uploading
achristes
Profile Blog Joined March 2011
Norway653 Posts
February 01 2012 15:36 GMT
#26
On February 02 2012 00:33 DarkPlasmaBall wrote:
Show nested quote +
On February 02 2012 00:31 achristes wrote:
On February 02 2012 00:24 DarkPlasmaBall wrote:
OP, your 2a is wrong as I explained before. Conditional probability doesn't work in that way because the two events are independent. The second event isn't conditioned on the first.

It's like saying, "I have a six-sided die and a coin. If I roll a 4 first, what's the chance of me flipping heads?" It's still 1/2. It doesn't change just because you roll the die in a certain way.

You didn't ask what the chance of Tor not going to a meeting *and* Nadia going to meeting is simultaneously (in that case, you would carry out the calculation the same way you did). The wording (particularly the "If") makes it a different question.

Nadia's probability of going to a meeting is independent of Tor, so whether or not Tor goes is irrelevant when deciding if Nadia goes. Therefore, Nadia's chances of going is still the established 70%. You don't need to include Tor's chance of not going, because- as was explicitly written in the instructions- the two events (Tor going and Nadia going) are independent events.

If it was slightly reworded as "What's the chance of Tor not going *and* Nadia going to the same meeting", then it would be .4 * .3 = .12.

Hope that helps

Oops, I'm so bad at writing these kinds of questions in english, thanks for telling me

But the bolded part should be .7 because it's the probability of Nadia going is 70%=0.7, or am I being stupid now?


I messed up the ending part to my previous post (should be ".4 * .7 = .28.") Edited it ^^

.4 for 1-.6 P(Tor not going)
.7 for P(Nadia going)
If it were worded differently.

It is now, mind checking for mistakes? ^^
youtube.com/spooderm4n | twitch.tv/spooderm4n | Random videos and games I feel like uploading
Koshi
Profile Blog Joined August 2010
Belgium38799 Posts
Last Edited: 2012-02-01 15:38:03
February 01 2012 15:36 GMT
#27
nvm. Going to read the blog. You people confuse me .

I had a good night of sleep.
JustPassingBy
Profile Blog Joined January 2011
10776 Posts
February 01 2012 15:43 GMT
#28
On February 01 2012 23:25 Plexa wrote:
Show nested quote +
On February 01 2012 23:21 ETisME wrote:
maths is fun until you get to a certain level where you starts to have questions that would get illogical (or basically you are required to abandon understanding the theory and learn to just DO maths), first one in my mind was sin, cos and tan.

If you like these kind of maths, stats are more to your taste to be honest

sin, cos and tan make perfect sense


Never wondered by those three are actually doing what they were meant to do?
(giving the ratio in a triangle with one 90° angle)

Well, back then, I did not. But now I am wondering why that question never came into my mind.
DarkPlasmaBall
Profile Blog Joined March 2010
United States46537 Posts
February 01 2012 15:43 GMT
#29
On February 02 2012 00:36 achristes wrote:
Show nested quote +
On February 02 2012 00:33 DarkPlasmaBall wrote:
On February 02 2012 00:31 achristes wrote:
On February 02 2012 00:24 DarkPlasmaBall wrote:
OP, your 2a is wrong as I explained before. Conditional probability doesn't work in that way because the two events are independent. The second event isn't conditioned on the first.

It's like saying, "I have a six-sided die and a coin. If I roll a 4 first, what's the chance of me flipping heads?" It's still 1/2. It doesn't change just because you roll the die in a certain way.

You didn't ask what the chance of Tor not going to a meeting *and* Nadia going to meeting is simultaneously (in that case, you would carry out the calculation the same way you did). The wording (particularly the "If") makes it a different question.

Nadia's probability of going to a meeting is independent of Tor, so whether or not Tor goes is irrelevant when deciding if Nadia goes. Therefore, Nadia's chances of going is still the established 70%. You don't need to include Tor's chance of not going, because- as was explicitly written in the instructions- the two events (Tor going and Nadia going) are independent events.

If it was slightly reworded as "What's the chance of Tor not going *and* Nadia going to the same meeting", then it would be .4 * .3 = .12.

Hope that helps

Oops, I'm so bad at writing these kinds of questions in english, thanks for telling me

But the bolded part should be .7 because it's the probability of Nadia going is 70%=0.7, or am I being stupid now?


I messed up the ending part to my previous post (should be ".4 * .7 = .28.") Edited it ^^

.4 for 1-.6 P(Tor not going)
.7 for P(Nadia going)
If it were worded differently.

It is now, mind checking for mistakes? ^^


If you want your answer to 2a to be correct (with the same explanation that you have written in your OP), I would recommend changing the question from

"If Tor doesn't go to one of the meetings, what is the probability of Nadia going to the same meeting?"

to

"What's the chance of Nadia going to a meeting that Tor doesn't go to?"

This explicitly shows that you need to multiply the probability that Tor doesn't go to a meeting with the probability that Nadia does go

Otherwise, it'll stay at 70% for the reasons explained before.
"There is nothing more satisfying than looking at a crowd of people and helping them get what I love." ~Day[9] Daily #100
achristes
Profile Blog Joined March 2011
Norway653 Posts
February 01 2012 15:45 GMT
#30
On February 02 2012 00:43 DarkPlasmaBall wrote:
Show nested quote +
On February 02 2012 00:36 achristes wrote:
On February 02 2012 00:33 DarkPlasmaBall wrote:
On February 02 2012 00:31 achristes wrote:
On February 02 2012 00:24 DarkPlasmaBall wrote:
OP, your 2a is wrong as I explained before. Conditional probability doesn't work in that way because the two events are independent. The second event isn't conditioned on the first.

It's like saying, "I have a six-sided die and a coin. If I roll a 4 first, what's the chance of me flipping heads?" It's still 1/2. It doesn't change just because you roll the die in a certain way.

You didn't ask what the chance of Tor not going to a meeting *and* Nadia going to meeting is simultaneously (in that case, you would carry out the calculation the same way you did). The wording (particularly the "If") makes it a different question.

Nadia's probability of going to a meeting is independent of Tor, so whether or not Tor goes is irrelevant when deciding if Nadia goes. Therefore, Nadia's chances of going is still the established 70%. You don't need to include Tor's chance of not going, because- as was explicitly written in the instructions- the two events (Tor going and Nadia going) are independent events.

If it was slightly reworded as "What's the chance of Tor not going *and* Nadia going to the same meeting", then it would be .4 * .3 = .12.

Hope that helps

Oops, I'm so bad at writing these kinds of questions in english, thanks for telling me

But the bolded part should be .7 because it's the probability of Nadia going is 70%=0.7, or am I being stupid now?


I messed up the ending part to my previous post (should be ".4 * .7 = .28.") Edited it ^^

.4 for 1-.6 P(Tor not going)
.7 for P(Nadia going)
If it were worded differently.

It is now, mind checking for mistakes? ^^


If you want your answer to 2a to be correct (with the same explanation that you have written in your OP), I would recommend changing the question from

"If Tor doesn't go to one of the meetings, what is the probability of Nadia going to the same meeting?"

to

"What's the chance of Nadia going to a meeting that Tor doesn't go to?"

This explicitly shows that you need to multiply the probability that Tor doesn't go to a meeting with the probability that Nadia does go

Otherwise, it'll stay at 70% for the reasons explained before.

Thank you, I will PM you if I decide to make another one
youtube.com/spooderm4n | twitch.tv/spooderm4n | Random videos and games I feel like uploading
Nehsb
Profile Joined May 2009
United States380 Posts
February 01 2012 15:46 GMT
#31
On February 02 2012 00:36 achristes wrote:
Show nested quote +
On February 02 2012 00:31 Treehead wrote:
On February 02 2012 00:03 achristes wrote:
On February 01 2012 23:39 Caller wrote:
This isn't math this is arithmetic

That statement makes no sense to me, because if you use numbers to find other number I call it math.


Mathematicians are constantly delving into supremely abstract realms to try to discover new theorems, corrollaries, etc. (mathematician's statements of truth) about extremely complicated situations. Some theorists work in areas that realistically may only be useful in practical ways to certain branches of physics. Mathematicians are creative thinkers, critical thinkers and problem solvers.

The OP is primarily concerned with questions of interpretation and execution of equations. Don't see the difference?

A problem for a mathematician goes something more like this:

The 'flea and comb space' is a topological space defined by a subspace of the two-dimensional coordinate plane which contains the point (0,1), all points (x,0), and all points (1/n,y), where n is any positive integer, and x and y are any real numbers between 0 and 1. Additionally, say that a point is "dense" (calling it "dense" because I actually don't remember the real term off hand) in the flea and comb set if it has points which are in the flea and comb set which are infinitely close to it, i.e. for any given arbitrarily small number, there is a point within that distance whcih is in the set. Prove that every point in this space is "dense" in the space, but that it is not the case that every point which is "dense" in the set is necessarily in the set.

The above is a relatively simple and straightforward math problem. More complicated problems may require a day or more of thought and reflection, experimentation and failure to resolve. This is the reason mathematicians bristle at the idea that saying 70% of $50 is $35 is "math" in the same sense. Math requires much more thought and effort.

So what you are saying is that mathematicians don't like that "normal" people say that the simpler parts of math is math?
I can understand it if that is the case though, but what am I supposed to call it then? ^^


That's why Caller's saying that it's arithmetic, not math.

@Treehead: I think the word you want is limit point?
DarkPlasmaBall
Profile Blog Joined March 2010
United States46537 Posts
Last Edited: 2012-02-01 16:03:39
February 01 2012 15:47 GMT
#32
On February 02 2012 00:45 achristes wrote:
Show nested quote +
On February 02 2012 00:43 DarkPlasmaBall wrote:
On February 02 2012 00:36 achristes wrote:
On February 02 2012 00:33 DarkPlasmaBall wrote:
On February 02 2012 00:31 achristes wrote:
On February 02 2012 00:24 DarkPlasmaBall wrote:
OP, your 2a is wrong as I explained before. Conditional probability doesn't work in that way because the two events are independent. The second event isn't conditioned on the first.

It's like saying, "I have a six-sided die and a coin. If I roll a 4 first, what's the chance of me flipping heads?" It's still 1/2. It doesn't change just because you roll the die in a certain way.

You didn't ask what the chance of Tor not going to a meeting *and* Nadia going to meeting is simultaneously (in that case, you would carry out the calculation the same way you did). The wording (particularly the "If") makes it a different question.

Nadia's probability of going to a meeting is independent of Tor, so whether or not Tor goes is irrelevant when deciding if Nadia goes. Therefore, Nadia's chances of going is still the established 70%. You don't need to include Tor's chance of not going, because- as was explicitly written in the instructions- the two events (Tor going and Nadia going) are independent events.

If it was slightly reworded as "What's the chance of Tor not going *and* Nadia going to the same meeting", then it would be .4 * .3 = .12.

Hope that helps

Oops, I'm so bad at writing these kinds of questions in english, thanks for telling me

But the bolded part should be .7 because it's the probability of Nadia going is 70%=0.7, or am I being stupid now?


I messed up the ending part to my previous post (should be ".4 * .7 = .28.") Edited it ^^

.4 for 1-.6 P(Tor not going)
.7 for P(Nadia going)
If it were worded differently.

It is now, mind checking for mistakes? ^^


If you want your answer to 2a to be correct (with the same explanation that you have written in your OP), I would recommend changing the question from

"If Tor doesn't go to one of the meetings, what is the probability of Nadia going to the same meeting?"

to

"What's the chance of Nadia going to a meeting that Tor doesn't go to?"

This explicitly shows that you need to multiply the probability that Tor doesn't go to a meeting with the probability that Nadia does go

Otherwise, it'll stay at 70% for the reasons explained before.

Thank you, I will PM you if I decide to make another one


Glad I could help The other ones look worded correctly ^^

EDIT: I haven't checked any other answers though.
"There is nothing more satisfying than looking at a crowd of people and helping them get what I love." ~Day[9] Daily #100
Koshi
Profile Blog Joined August 2010
Belgium38799 Posts
February 01 2012 15:49 GMT
#33
2d seems wrong to me. You probably didnt fill in the complete answer or something.

Both going to 3 consecutive meeting would be
(0,6*0,7) ^ 3 = 7,4%
I had a good night of sleep.
Iranon
Profile Blog Joined March 2010
United States983 Posts
February 01 2012 15:49 GMT
#34
On February 02 2012 00:36 achristes wrote:
Show nested quote +
On February 02 2012 00:31 Treehead wrote:
On February 02 2012 00:03 achristes wrote:
On February 01 2012 23:39 Caller wrote:
This isn't math this is arithmetic

That statement makes no sense to me, because if you use numbers to find other number I call it math.


Mathematicians are constantly delving into supremely abstract realms to try to discover new theorems, corrollaries, etc. (mathematician's statements of truth) about extremely complicated situations. Some theorists work in areas that realistically may only be useful in practical ways to certain branches of physics. Mathematicians are creative thinkers, critical thinkers and problem solvers.

The OP is primarily concerned with questions of interpretation and execution of equations. Don't see the difference?

A problem for a mathematician goes something more like this:

The 'flea and comb space' is a topological space defined by a subspace of the two-dimensional coordinate plane which contains the point (0,1), all points (x,0), and all points (1/n,y), where n is any positive integer, and x and y are any real numbers between 0 and 1. Additionally, say that a point is "dense" (calling it "dense" because I actually don't remember the real term off hand) in the flea and comb set if it has points which are in the flea and comb set which are infinitely close to it, i.e. for any given arbitrarily small number, there is a point within that distance whcih is in the set. Prove that every point in this space is "dense" in the space, but that it is not the case that every point which is "dense" in the set is necessarily in the set.

The above is a relatively simple and straightforward math problem. More complicated problems may require a day or more of thought and reflection, experimentation and failure to resolve. This is the reason mathematicians bristle at the idea that saying 70% of $50 is $35 is "math" in the same sense. Math requires much more thought and effort.

So what you are saying is that mathematicians don't like that "normal" people say that the simpler parts of math is math?
I can understand it if that is the case though, but what am I supposed to call it then? ^^


No. What he's saying is that mathematicians know what math is, and most people think that what they learned in high school is basically all there is to mathematics, with some additional complexity tacked on.

Addition and division and calculus and all the other parts of math that are routine calculations are lumped together as arithmetic because they are to math as paint-by-numbers is to watercolor painting. Sure, it's the same thing in a technical sense -- you have some brushes, and some water, and some paint, and you put paint and water on the brushes and make parts of the paper turn different colors, but it's not the same thing in a meaningful sense, as one is Art and the other is decidely not. THAT'S what people mean when they say that something which is clearly a mathematical statement or problem is "not real math". The beginnings of math is just about calculating some quantity. That's what people mean by arithmetic. The rest of math bears some resemblance to arithmetic, in that it involves formal symbolic manipulations to draw conclusions about the properties of certain abstract objects based on properties of other related objects, but that's about as far as the resemblance goes. (Unless you're in applied math, in which case I'm not talking to you).
Koshi
Profile Blog Joined August 2010
Belgium38799 Posts
Last Edited: 2012-02-01 15:58:43
February 01 2012 15:57 GMT
#35
3c is also wrong.

10.000 + 120 * x = 200 * x (x = the amount of attendees)
x = 50 + 0,6 x
0,4 x = 50
x = 125

So the answer is 125 students


EDIT: I am going home now, I might check the other ones when I get home .
I had a good night of sleep.
.Sic.
Profile Joined February 2011
Korea (South)497 Posts
February 01 2012 15:57 GMT
#36
On February 01 2012 23:21 ETisME wrote:
maths is fun until you get to a certain level where you starts to have questions that would get illogical (or basically you are required to abandon understanding the theory and learn to just DO maths), first one in my mind was sin, cos and tan.

If you like these kind of maths, stats are more to your taste to be honest


..... math is always pure logic. If you're getting to a point where math seems illogical, that means you don't have a logical understanding of the concepts.
Clan MvP Member | http://sc2ranks.com/kr/3273340/SicMvP
DarkPlasmaBall
Profile Blog Joined March 2010
United States46537 Posts
Last Edited: 2012-02-01 16:05:23
February 01 2012 16:03 GMT
#37
On February 02 2012 00:49 Koshi wrote:
2d seems wrong to me. You probably didnt fill in the complete answer or something.

Both going to 3 consecutive meeting would be
(0,6*0,7) ^ 3 = 7,4%


I agree with this.

P(Nadia going) * P(Tor going) = P(Nadia and Tor going to the same)

to the third power, for three meetings.
"There is nothing more satisfying than looking at a crowd of people and helping them get what I love." ~Day[9] Daily #100
achristes
Profile Blog Joined March 2011
Norway653 Posts
February 01 2012 16:06 GMT
#38
On February 02 2012 01:03 DarkPlasmaBall wrote:
Show nested quote +
On February 02 2012 00:49 Koshi wrote:
2d seems wrong to me. You probably didnt fill in the complete answer or something.

Both going to 3 consecutive meeting would be
(0,6*0,7) ^ 3 = 7,4%


I agree with this.

P(Nadia going) * P(Tor going) = P(Nadia and Tor going to the same)

to the third power, for three meetings.

Sorry, my bad.
youtube.com/spooderm4n | twitch.tv/spooderm4n | Random videos and games I feel like uploading
DarkPlasmaBall
Profile Blog Joined March 2010
United States46537 Posts
February 01 2012 16:09 GMT
#39
On February 02 2012 01:06 achristes wrote:
Show nested quote +
On February 02 2012 01:03 DarkPlasmaBall wrote:
On February 02 2012 00:49 Koshi wrote:
2d seems wrong to me. You probably didnt fill in the complete answer or something.

Both going to 3 consecutive meeting would be
(0,6*0,7) ^ 3 = 7,4%


I agree with this.

P(Nadia going) * P(Tor going) = P(Nadia and Tor going to the same)

to the third power, for three meetings.

Sorry, my bad.


No worries. Are these questions you got wrong on a test? Or questions you got correct but just mis-translated into English? Or something else? Your OP says they were on your last math test, so are you looking to see if you got the answers right? Or are testing us?
"There is nothing more satisfying than looking at a crowd of people and helping them get what I love." ~Day[9] Daily #100
Treehead
Profile Blog Joined November 2010
999 Posts
Last Edited: 2012-02-01 16:16:35
February 01 2012 16:12 GMT
#40
On February 02 2012 00:36 achristes wrote:
Show nested quote +
On February 02 2012 00:31 Treehead wrote:
On February 02 2012 00:03 achristes wrote:
On February 01 2012 23:39 Caller wrote:
This isn't math this is arithmetic

That statement makes no sense to me, because if you use numbers to find other number I call it math.


Mathematicians are constantly delving into supremely abstract realms to try to discover new theorems, corrollaries, etc. (mathematician's statements of truth) about extremely complicated situations. Some theorists work in areas that realistically may only be useful in practical ways to certain branches of physics. Mathematicians are creative thinkers, critical thinkers and problem solvers.

The OP is primarily concerned with questions of interpretation and execution of equations. Don't see the difference?

A problem for a mathematician goes something more like this:

The 'flea and comb space' is a topological space defined by a subspace of the two-dimensional coordinate plane which contains the point (0,1), all points (x,0), and all points (1/n,y), where n is any positive integer, and x and y are any real numbers between 0 and 1. Additionally, say that a point is "dense" (calling it "dense" because I actually don't remember the real term off hand) in the flea and comb set if it has points which are in the flea and comb set which are infinitely close to it, i.e. for any given arbitrarily small number, there is a point within that distance whcih is in the set. Prove that every point in this space is "dense" in the space, but that it is not the case that every point which is "dense" in the set is necessarily in the set.

The above is a relatively simple and straightforward math problem. More complicated problems may require a day or more of thought and reflection, experimentation and failure to resolve. This is the reason mathematicians bristle at the idea that saying 70% of $50 is $35 is "math" in the same sense. Math requires much more thought and effort.

So what you are saying is that mathematicians don't like that "normal" people say that the simpler parts of math is math?
I can understand it if that is the case though, but what am I supposed to call it then? ^^


Let me put it in SC2 terms. Let's say I have a simulator which I have programmed to simulate an SC2 game for one player with clear goals in mind - the computer makes building choices based on a priority list I determined for it before the game. We assume a certain rush distance and a few other assumptions, but there's no map, no micro and the units just kind of attack each other simulatedly when the decision to attack comes up.

I wrote the priority list. I told it what it would need to attack. The units are SC2 units. Am I playing SC2? I could say that I'm playing SC2 but making it easier, but you'd tell me I missed the point of the game completely (I hope).

Math is about being creative, then being critical, then putting your work out on paper, the last step of which is the easiest, and the only one involved in arithmetic - that you write the numbers on the paper in the right order and remember how they go together. Nothing is being created, nothing is being critiqued, nothing is being improvised - it's rote. Math is far from being rote.

On February 02 2012 00:46 Nehsb wrote:
Show nested quote +
On February 02 2012 00:36 achristes wrote:
On February 02 2012 00:31 Treehead wrote:
On February 02 2012 00:03 achristes wrote:
On February 01 2012 23:39 Caller wrote:
This isn't math this is arithmetic

That statement makes no sense to me, because if you use numbers to find other number I call it math.


Mathematicians are constantly delving into supremely abstract realms to try to discover new theorems, corrollaries, etc. (mathematician's statements of truth) about extremely complicated situations. Some theorists work in areas that realistically may only be useful in practical ways to certain branches of physics. Mathematicians are creative thinkers, critical thinkers and problem solvers.

The OP is primarily concerned with questions of interpretation and execution of equations. Don't see the difference?

A problem for a mathematician goes something more like this:

The 'flea and comb space' is a topological space defined by a subspace of the two-dimensional coordinate plane which contains the point (0,1), all points (x,0), and all points (1/n,y), where n is any positive integer, and x and y are any real numbers between 0 and 1. Additionally, say that a point is "dense" (calling it "dense" because I actually don't remember the real term off hand) in the flea and comb set if it has points which are in the flea and comb set which are infinitely close to it, i.e. for any given arbitrarily small number, there is a point within that distance whcih is in the set. Prove that every point in this space is "dense" in the space, but that it is not the case that every point which is "dense" in the set is necessarily in the set.

The above is a relatively simple and straightforward math problem. More complicated problems may require a day or more of thought and reflection, experimentation and failure to resolve. This is the reason mathematicians bristle at the idea that saying 70% of $50 is $35 is "math" in the same sense. Math requires much more thought and effort.

So what you are saying is that mathematicians don't like that "normal" people say that the simpler parts of math is math?
I can understand it if that is the case though, but what am I supposed to call it then? ^^


That's why Caller's saying that it's arithmetic, not math.

@Treehead: I think the word you want is limit point?


Yeah, I tried to google "sequence of points approaching" and it gave me "limit", but somehow I thought there was another word for it in topology. Like a set which is closed contains its'... "boundary"? Maybe that's the right word. Idk - it's been at least 5 years since I did topology, and honestly I'm not sure why that's the problem that came to mind when I picked one out of my graduate career.
Prev 1 2 3 Next All
Please log in or register to reply.
Live Events Refresh
Next event in 6m
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
deth 31
StarCraft: Brood War
Rain 3898
FanTaSy 770
BeSt 463
Hyuk 404
Larva 392
EffOrt 281
Stork 276
Hyun 264
actioN 236
firebathero 228
[ Show more ]
Zeus 165
ZerO 143
Barracks 102
Mind 79
ggaemo 67
Sharp 47
ToSsGirL 40
Shinee 31
Aegong 28
sorry 24
Free 22
soO 20
HiyA 18
ZergMaN 16
910 14
Sacsri 12
scan(afreeca) 11
yabsab 10
Noble 9
ajuk12(nOOB) 7
Dota 2
ODPixel171
League of Legends
Reynor94
Other Games
gofns42514
singsing1801
Dewaltoss21
BEARDiaguz10
ZerO(Twitch)6
Pyrionflax2
Organizations
Other Games
gamesdonequick575
[ Show 10 non-featured ]
StarCraft 2
• LUISG 35
• Dystopia_ 3
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Migwel
StarCraft: Brood War
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
Upcoming Events
CranKy Ducklings
6m
WardiTV Invitational
1h 6m
ByuN vs Percival
Percival vs SHIN
Percival vs herO
ByuN vs SHIN
ByuN vs herO
SHIN vs herO
AI Arena Tournament
6h 6m
BSL Open Qualifier
9h 6m
Sparkling Tuna Cup
1d
OSC
1d 3h
IPSL
1d 6h
BSL Open Qualifier
1d 9h
WardiTV Weekly
2 days
IntoTheTV X SOOP
3 days
[ Show More ]
PiGosaur Cup
3 days
The PondCast
4 days
Kung Fu Cup
4 days
IntoTheTV X SOOP
5 days
FEL
6 days
Liquipedia Results

Completed

Super Anchor Qualifying S3
Blizzard Classic Cup 2026
Copium Cup

Ongoing

ASL Season 22
Acropolis #5
KCM Race Survival 2026 Season 4
Acropolis #5 - GSC
HCC Season 3
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026
Esports World Cup 2026

Upcoming

BSL Season 23
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
IEM Beijing 2026
Stake Ranked Episode 5
PGL Masters Bucharest 2026
1win Private Club #2
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.